OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) − 4

OpenStudy (anonymous):

3x +3 =-2x+2-4 5x=-2-3=-5 x=-1

OpenStudy (anonymous):

WHAT YOU DO IS YOU PUT ALL THE VARIABLES ON ONE SIDE {THE MOST INPORTENT RULE YOU NEED TO REMEMBER IS THAT THE EQUAL IS THE SPLITTING LINE